all kings and queens are removed from  a pack of 52 cards the remaining cards are well shuffled and then a card is randomly from it. find the probability that this card is 
a- a red face card
b-a black card

Grade:10

4 Answers

Piyush Kumar Behera
7 years ago
See when you remove all king and queen ,4 red face card and 4 black face card is removed.
Total cards remaining=52-8=44
red face card remaining=6-4=2
as 4 black card are removed so 26-4=22 cards are remaining
the probability that this card is
a- a red face card=1/22
b-a black card=1/2
Please approve this solution to encourage me
Hinal Srivastava
7 years ago
We know that there are 4 kings and 4 queens in a deck
Hence, the remaining number of cards= 52-(4+4)=44
 
a) Number of red face cards=2 i.e  2 jacks
      PROBABILITY= 2/44=1/22
  1. ​Number of black cards=26-4=22
            PROBABILITY= 22/44=1/2
           HOPE THIS SATISFIES YOU!!
sachin
6 years ago
 
Total cards remaining=52-8=44
red face card remaining=6-4=2
as 4 black card are removed so 26-4=22 cards are remaining
the probability that this card is
a- a red face card=1/22
b-a black card=1/2
sachin
6 years ago
 
Total cards remaining=52-8=44
red face card remaining=6-4=2
as 4 black card are removed so 26-4=22 cards are remaining
the probability that this card is
a- a red face card=1/22
b-a black card=1/2.